OpenStudy (anonymous):

help! find the solutions x^4+25x^2+144=0

OpenStudy (anonymous):

Factor.

OpenStudy (jdoe0001):

$$ x^4+\color{blue}{25^2}+144=0?\\ \text{are you sure it isn't}\\ x^4+\color{blue}{24^2}+144=0 $$

OpenStudy (jdoe0001):

$$ x^4+\color{blue}{25x^2}+144=0?\\ \text{are you sure it isn't}\\ x^4+\color{blue}{24x^2}+144=0 $$ rather :/

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

your equation, that is :)

OpenStudy (anonymous):

It doesn't have to be. There's a factorization that works.

OpenStudy (jdoe0001):

well, true, doesn't have to be a perfect trinomial, ok

OpenStudy (jdoe0001):

just wondering :)

OpenStudy (anonymous):

It's actually much better the way it factors, you don't end up with sqrt(12)